Three point charges are arranged in a horizontal line as shown below. Find the electric forces (in units of kQ2/R2) on the charges given the following: Q1 = -9 Q, Q2 = -125 Q, Q3 = -64 Q, r1 = 3 R, and r2 = 5 R. Remember that a positive force points to the right and a negative force points to the left. What is the force on charge Q1 due to Q2? kQ2/R2 Tries 0/2 What is the force on charge Q1 due to Q3? kQ2/R2 Tries 0/2 What is the force on charge Q2 due to Q3? kQ2/R2 Tries 0/2 What is the force on charge Q2 due to Q1? kQ2/R2 Tries 0/2 What is the force on charge Q3 due to Q1? kQ2/R2 Tries 0/2 What is the force on charge Q3 due to Q2? kQ2/R2 Tries 0/2 What is the net force on charge Q1? kQ2/R2 Tries 0/2 What is the net force on charge Q2? kQ2/R2 Tries 0/2 What is the net force on charge Q3? kQ2/R2 Tries 0/2 What is the sum of the forces on all three charges? kQ2/R2
Answer #1

1) F12 = -k(9Q)(125Q)/9R2 (to the left) = -125 KQ2/R2 (repulsive force to the left)
answer = -125

2) F13 = -k(9Q)(64Q)/64R2 (to the left) = -9 KQ2/R2 (repulsive force to the left)
answer = -9

3) F23 = -k(125Q)(64Q)/25R2 (to the left) = -320 KQ2/R2 (repulsive force to the left)
answer = -320

4) F21 = = k(125Q)(9Q)/9R2 (to the right) = 125 KQ2/R2 (repulsive force to the right)
answer = 125

5) F31 = = k(9Q)(64Q)/64R2 (to the right) = 9 KQ2/R2 (repulsive force to the right)
answer = 9

6) F32 = k(125Q)(64Q)/25R2 (to the right) = 320 KQ2/R2 (repulsive force to the right)
answer = 320

7) Net force on Q1, F1 = F12 + F13 = -134

8) Net force on Q2, F2 = F23 + F21 = -195

9) Net force on Q3, F1 = F31 + F32 = 329

10) Sum of all forces = F1 +F2 +F3 = -134-195+329 = 0